An aerosol can contains gases under a pressure of 4.5 atm at 27 â—¦C. If the can is left on a hot sandy beach, the pressure of the gases increases to 4.85 atm. What is the Celsius temperature on the beach?


1
Expert's answer
2021-10-18T06:49:27-0400

P1 = 4.50 atm 

T1 =  27◦C +273 = 300 K

V1 = V2 = const

P2 = 4.85 atm

T2 = unknown


Solution:

Since the volume doesn't change, Gay-Lussac's Law can be used.

Gay-Lussac's Law can be expressed as: P1/T1 = P2/T2

P1T2 = P2T1

To find the final temperature, solve the equation for T2:

T2 = P2T1/P1

T2 = (4.85 atm × 300 K) / (4.50 atm) = 323.33 K = 323 K

T2 = 323 K or 323 - 273 = 50◦C


Answer: 50â—¦C